# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

### v3

@@ III-D $\epsilon-\chi$ Tradeoff in the AWGN Channel
In this subsection, we will take AWGN as an example to characterize the $\epsilon-\chi$ tradeoff through an inequality. The AWGN channel model is widely used in wired communications and satellite communications. With the rapid development of deep learning technology, the communication between processing units, which usually belongs to wired communication, has also raised much attention from both academia and industry. Thus, characterizing the performance limit of the AWGN channel in the FBL regime is important. Different from the discussion in Subsections III-A, III-B, and III-C, we do not use the assumptions shown in Theorem 1 in this subsection.
For the AWGN channel, the DVP is equivalent to QVP since the service rate is fixed given $N$, $\epsilon$, and $\gamma$. Thus, we can choose to use QVP to represent the delay performance of the communication system in the AWGN channel. Similarly, we use $\epsilon$ to evaluate the reliability of the system. Based on the above considerations, we obtain the $\epsilon-\chi$ tradeoff in an inequality in Corollary 3. With some abuse of notations, we let $C=\log_{2}(1+\gamma)$ and $V=\log_{2}^{2}e\left(1-\frac{1}{(1+\gamma)^{2}}\right)$ in Corollary 3.
@@
${\color[rgb]{0,0,0}\epsilon\geq Q\left(\frac{NC+\frac{\log_{2}N}{2}+G(N,\gamma,\epsilon)-\alpha_{A}\left(-\frac{1}{L}\ln\chi_{\rm th}\right)}{\sqrt{NV}}\right).}$ (70)
@@
See Appendix G.
Corollary 3 presents the relationships between $\epsilon$ and $\chi$. In practical applications, we can use these relationships to determine the proper choice of $\epsilon$ with the given $\chi_{\rm th}$. For the fading channel, we can conduct a similar analysis. However, it is not trivial because it is hard to obtain the analytical expression of $\theta$. With the help of approximation methods and under some assumptions, one may obtain the analytical or even closed-form approximation of $\theta$ [45]. Subsequently, a similar expression to Eq. (70) can be derived, which is beyond the scope of this paper.
@@ IV Effective Capacity Approximation with Normal Approximation
In this section, we present an EC approximation method for low-latency communication systems. The closed-form expression of the approximation characterizes the influence of the channel distribution, error probability, and blocklength on the EC. As shown in Eqs. (19) and (21), there are limitations on the use of LDT to determine the QVP and DVP in URLLC scenarios. However, EC is widely used in the works for URLLC scenarios [26, 27, 24, 25] to evaluate the QoS-constrained throughput. In these studies, the value of the QoS exponent $\theta$ is given to represent the QoS requirement, while the EC is adopted to characterize the throughput under the QoS requirement. Some scheduling schemes in URLLC were proposed to maximize the EC in these works. Therefore, deriving the closed-form expression of EC is crucial for further optimization. However, it is challenging to derive a closed-form expression of the EC, complicating throughput analysis and the comparison between different scheduling schemes. To address this issue, we propose a Laplace’s-method-based approximation for EC in this section. In this subsection, we set $T\geq 1$, which covers the general low-latency scenarios. Note that we use normal approximation to approximate the maximum coding rate in this section. Thus, the proposed approximation can serve as an efficient EC approximation method in scenarios where normal approximation achieves satisfactory accuracy.
@@ IV-A EC Approximation for Fixed-Power Allocation
In this subsection, we start with the simple case in which the transmitted power is constant in each time slot. For low-latency scenarios, the threshold $L$ and $D_{\rm max}$ are usually small due to the stringent QoS requirement [22]. Thus, to achieve an extremely small QVP or DVP, the value of the QoS exponent $\theta$ or the revised QoS exponent $\theta N$ needs to be large. By this means, $\theta NT$ should have a large value. Based on this thought, we start our analysis of EC with the fixed-power allocation in Lemma 2. Note that $\ell=\left(\frac{Q^{-1}(\epsilon)}{\sqrt{N}}\right)^{2}$ as defined before Theorem 4.
Lemma 2. For a low-latency communication system adopting fixed-power allocation, EC is approximated as
$\alpha_{S}(\theta)=Nr(x^{*})-\frac{\ln f(x^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right),$ (71)
where $r(x)$ is defined in Eq. (40) and $x^{*}$ is given by
@@
See Appendix H.
In Lemma 2, we provide an approximation of EC for the fixed-power allocation, which provides a closed-form expression of EC with arbitrary fading channels. Besides, the approximation error is $O\left(\frac{1}{\theta^{2}T^{2}N}\right)$. Thus, as we mentioned before Lemma 2, this approximation performs better in a low-latency communication system, which we will validate by numerical calculation in Section V.
@@ IV-B EC Approximation with Channel-Gain-Based Power Allocation
In this subsection, we will generalize the proposed EC approximation method for the fixed-power allocation scheme to certain types of channel-gain-based power allocation schemes. The conditions for using the proposed approximation are presented. Moreover, we discuss the applicability of the proposed approximation for some classical channel-gain-based power allocation schemes.
Based on Lemma 2, we can extend the approximation to more general power allocation schemes. We summarize this generalization in Theorem 8.
Theorem 8. For a low-latency communication system which adopts a channel-gain-based power allocation scheme $\Xi(x)$ satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ with channel distribution $f(\cdot)$, EC is approximated as
$\alpha_{S}(\theta)=Nr(x^{*})-\frac{\ln f(x^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right),$ (73)
where $r(x)$ is defined in Eq. (40) and $x^{*}$ is given by
$\Xi(x^{*})x^{*}=\sqrt{\frac{1+\sqrt{1+4\ell}}{2}}-1.$ (74)
@@
See Appendix I.
In Theorem 8, we propose the EC approximation for a specific kind of power allocation scheme. For the policies satisfying $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$, $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ holds. $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$ means that the allocated power does not decrease with the increasing channel power gain. This condition can be met with some power allocation schemes, one of which is the well-known water-filling scheme [46]. Furthermore, when $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$, $\Xi(x)x$ is a monotonic function of $x$. Thus, the solution of Eq. (42) can be obtained directly via binary search.
A power-control policy that maximizes the effective capacity under the average power constraint is proposed in [47]. This policy also satisfies the constraint presented in Theorem 8. Let $\Xi_{1}(x)$ denote the power-control policy proposed in [47], which can be expressed as
@@
where $a_{1}$ and $a_{2}$ are positive constants. $a_{1}$ is related to the average power constraint, while $a_{2}$ is a linear and increasing function of $\theta$. For $x\geq\frac{a_{1}}{\gamma}$, we have
$\Xi_{1}(x)+\frac{{\rm d}\Xi_{1}(x)}{{\rm d}x}x$ $=\frac{1}{a_{1}^{\frac{1}{a_{2}+1}}(\gamma x)^{\frac{a_{2}}{a_{2}+1}}}-\frac{1}{\gamma x}+x\left(-\frac{a_{2}}{a_{2}+1}\frac{1}{a_{1}^{\frac{1}{a_{2}+1}}\gamma^{\frac{a_{2}}{a_{2}+1}}}x^{-\frac{a_{2}}{a_{2}+1}-1}+\frac{1}{\gamma}\frac{1}{x^{2}}\right)$ (76)
$=\frac{1-\frac{a_{2}}{a_{2}+1}}{a_{1}^{\frac{1}{a_{2}+1}}\gamma^{\frac{a_{2}}{a_{2}+1}}}x^{-\frac{a_{2}}{a_{2}+1}}.$
Since $\frac{a_{2}}{a_{2}+1}<1$, it follows that $\Xi_{1}(x)+\frac{{\rm d}\Xi_{1}(x)}{{\rm d}x}x\geq 0$ for $x>0$, which means the power-control policy proposed in [47] satisfies the condition of Theorem 8. There are two special cases to consider. When $a_{2}$ approaches 0, the power-control policy proposed in [47] degenerates to the water-filling policy, which has been discussed above. When $a_{2}$ approaches infinity, the power-control policy proposed in [47] degenerates to channel inversion policy, with which the rate is constant over all channel states. In this case, EC is trivial, as the rate remains fixed.
Although the approximation provided in Theorem 8 can be applied to the policies satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$, the accuracy of this approximation can be further enhanced through additional optimization tailored to the specific properties of each policy. For example, in the water-filling scheme, there is a channel-gain threshold $h_{\rm th}$. If the channel gain falls below this threshold, the allocated power is 0, which means the transmission rate is constant in the FBL regime when $h[n]<h_{\rm th}$. [^9] The true rate should be zero when $\gamma=0$. However, since we use the normal approximation as we mentioned at the beginning of this section, we let the rate be equal to $\frac{\log_{2}N}{2}$ when $\gamma=0$ according to (9). This setting can be easily adjusted by setting the rate to 0 when $\gamma=0$, which does not affect the accuracy of this approximation. Let $\xi(\theta)=e^{-\theta T\frac{\log_{2}(N)}{2}}\int_{0}^{h_{\rm th}}f(x){\rm d}x$. Based on the above analysis and Theorem 8, EC can be approximated as
For those policies exhibiting truncated characteristics, Eq. (77) provides a more accurate approximation of their EC.
For the low-latency system with the retransmission mechanism, the EC is discussed in [28], which differs from the above analysis without the retransmission mechanism. In these works, a simple ARQ mechanism is assumed to be adopted. With this mechanism, the receiver sends a negative acknowledgment requesting the retransmission of the message in case of an erroneous reception. In Corollary 4, we generalize the EC approximation to the low-latency communication system incorporating this ARQ mechanism.
Corollary 4. For a low-latency system with the above ARQ mechanism and a channel-gain-based power allocation scheme $\Xi(x)$ satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ with channel distribution $f(\cdot)$, EC is approximated as
$\alpha_{S}(\theta)=-\frac{1}{\theta T}\ln\left(\epsilon+(1-\epsilon)\sqrt{\frac{2\pi}{\theta NT|r^{\prime\prime}(x^{*})|}}f(x^{*})e^{-\theta NTr(x^{*})}\left(1+O\left((\theta NT)^{-1}\right)\right)\right),$ (78)
where $r(x)$ is defined in Eq. (40) and $x^{*}$ is given by Eq. (42).
Proof:
See Appendix J.

### v4

@@ III-D $\epsilon-\chi$ Tradeoff in the AWGN Channel
In this subsection, we will take AWGN as an example to characterize the $\epsilon-\chi$ tradeoff through an inequality. The AWGN channel model is widely employed in wired communications and satellite communications. With the rapid development of deep learning technologies, communication between processing units, which usually belongs to wired communication, has also raised much attention from both academia and industry. Therefore, characterizing the performance limits of the AWGN channel in the FBL regime is important. Different from the discussion in Subsections III-A, III-B, and III-C, we will not adopt the assumptions presented in Theorem 1 in this subsection.
For the AWGN channel, the DVP is equivalent to QVP since the service rate is fixed given $N$, $\epsilon$, and $\gamma$. Therefore, we can choose to use QVP representing the delay performance of the communication system in the AWGN channel. Similarly, we use $\epsilon$ to evaluate the reliability of the system. Based on these considerations, we derive the $\epsilon-\chi$ tradeoff as an inequality in Corollary 3. With some abuse of notations, we let $C=\log_{2}(1+\gamma)$ and $V=\log_{2}^{2}e\left(1-\frac{1}{(1+\gamma)^{2}}\right)$ in Corollary 3.
@@
$\epsilon\geq Q\left(\frac{NC+\frac{\log_{2}N}{2}+G(N,\gamma,\epsilon)-\alpha_{A}\left(-\frac{1}{L}\ln\chi_{\rm th}\right)}{\sqrt{NV}}\right).$ (71)
@@
See Appendix G. ∎
Corollary 3 presents the relationships between $\epsilon$ and $\chi$. In practical applications, we can use these relationships to determine the appropriate choice of $\epsilon$ for a given $\chi_{\rm th}$. For the fading channel, a similar analysis can be performed. However, it is not trivial, as obtaining the analytical expression for $\theta$ is challenging. With the help of approximation methods and under certain assumptions, one may derive an analytical or even closed-form approximation for $\theta$ [45]. Subsequently, a similar expression to Eq. (71) can be derived, which is beyond the scope of this paper.
$\alpha_{S}(\theta)=Nr(x_{1}^{*})-\frac{\ln f(x_{1}^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x_{1}^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right).$ (72)
$\alpha_{S}(\theta)=Nr(x_{2}^{*})-\frac{\ln f(x_{2}^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x_{2}^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right).$ (74)
$\Xi_{1}(x)+\frac{{\rm d}\Xi_{1}(x)}{{\rm d}x}x$ $=\frac{1}{a_{1}^{\frac{1}{a_{2}+1}}(\gamma x)^{\frac{a_{2}}{a_{2}+1}}}-\frac{1}{\gamma x}+x\left(-\frac{a_{2}}{a_{2}+1}\frac{1}{a_{1}^{\frac{1}{a_{2}+1}}\gamma^{\frac{a_{2}}{a_{2}+1}}}x^{-\frac{a_{2}}{a_{2}+1}-1}+\frac{1}{\gamma}\frac{1}{x^{2}}\right)=\frac{1-\frac{a_{2}}{a_{2}+1}}{a_{1}^{\frac{1}{a_{2}+1}}\gamma^{\frac{a_{2}}{a_{2}+1}}}x^{-\frac{a_{2}}{a_{2}+1}}.$ (77)
@@ IV Effective Capacity Approximation with Normal Approximation
In this section, we present an EC approximation method for low-latency communication systems. The closed-form expression of this approximation characterizes the influence of the channel distribution, error probability, and blocklength on EC. As shown in Eqs. (19) and (21), there are limitations in using LDT to determine the QVP and DVP in URLLC scenarios. However, EC has been widely adopted in various studies of URLLC to evaluate the QoS-constrained throughput [26, 27, 24, 25]. In these studies, the value of the QoS exponent $\theta$ is typically given to represent the QoS requirement, while EC is employed to quantify the throughput subject to these QoS requirements. Several scheduling schemes for URLLC have been proposed to maximize EC in these works. Therefore, deriving a closed-form expression of EC is crucial for further system optimization. However, it is challenging to derive such a closed-form expression for EC, which complicates throughput analysis and the comparison of different scheduling schemes. To address this issue, we propose a Laplace’s-method-based approximation for EC in this section. In this subsection, we set $T\geq 1$, which covers the general low-latency scenarios. Note that we use normal approximation to approximate the maximum coding rate in this section. Thus, the proposed approximation can serve as an efficient EC approximation method for scenarios where normal approximation achieves satisfactory accuracy.
@@ IV-A EC Approximation for Fixed-Power Allocation
In this subsection, we start with a simple case in which the transmitted power remains constant in each time slot. For low-latency scenarios, the thresholds for queue length $L$ and latency $D_{\rm max}$ are usually small due to the stringent QoS requirements [22]. Thus, to achieve extremely small QVP or DVP, the value of the QoS exponent $\theta$ or the revised QoS exponent $\theta N$ needs to be large. Consequently, $\theta NT$ should have a large value. With this consideration, we begin our analysis by the fixed-power allocation in Lemma 2. Note that $\ell=\left(\frac{Q^{-1}(\epsilon)}{\sqrt{N}}\right)^{2}$, as defined before Theorem 4.
Lemma 2. For a low-latency communication system adopting fixed-power allocation, EC is approximated using Eq. (72), which is shown at the bottom of this page. In Eq. (72), $r(x)$ is defined in Eq. (41), and $x_{1}^{*}$ is given by
$x_{1}^{*}=\frac{1}{\gamma}\left(\sqrt{\frac{1+\sqrt{1+4\ell}}{2}}-1\right).$ (73)
@@
See Appendix H. ∎
In Lemma 2, we provide an approximation of EC for the fixed-power allocation, which provides a closed-form expression of EC applicable to arbitrary fading channels. Additionally, the approximation error is $O\left(\frac{1}{\theta^{2}T^{2}N}\right)$. Thus, as mentioned earlier, this approximation performs better in a low-latency communication system, which we will validate through numerical calculation in Section V.
@@ IV-B EC Approximation with Channel-Gain-Based Power Allocation
In this subsection, we will generalize the proposed EC approximation method for the fixed-power allocation scheme to certain types of channel-gain-based power allocation schemes. The conditions for using the proposed approximation are provided. Moreover, we discuss the applicability of the proposed approximation for several classical channel-gain-based power allocation schemes.
Building on Lemma 2, we extend the approximation to more general power allocation schemes, which we summarize in Theorem 8.
Theorem 8. For a low-latency communication system which adopts a channel-gain-based power allocation scheme $\Xi(x)$ satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ with a channel distribution $f(\cdot)$, EC is approximated using Eq. (74), which is presented at the bottom of this page. In Eq. (74), $r(x)$ is defined in Eq. (41), and $x_{2}^{*}$ is given by
$\Xi(x_{2}^{*})x_{2}^{*}=\sqrt{\frac{1+\sqrt{1+4\ell}}{2}}-1.$ (75)
@@
See Appendix I. ∎
In Theorem 8, we present an EC approximation for a specific kind of power allocation schemes. Specifically, for the policies satisfying $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$, the condition $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ is satisfied. The requirement $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$ indicates that the allocated power does not decrease as the channel power gain increases. This condition holds for certain power allocation schemes, including the well-established water-filling scheme [46]. Furthermore, when $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$, $\Xi(x)x$ is a monotonic function of $x$. Consequently, the solution to Eq. (43) can be efficiently obtained via the binary search algorithm.
A power-control policy designed to maximize EC while adhering to the average power constraint is proposed in [16]. This policy also complies with the constraints outlined in Theorem 8. Let $\Xi_{1}(x)$ denote the power-control policy introduced in [16], which can be expressed as
@@
where $a_{1}$ and $a_{2}$ are positive constants. $a_{1}$ is related to the average power constraint, while $a_{2}$ is a linear and increasing function of $\theta$. For $x\geq\frac{a_{1}}{\gamma}$, we have Eq. (77), which is provided at the bottom of this page. Since $\frac{a_{2}}{a_{2}+1}<1$, it follows that $\Xi_{1}(x)+\frac{{\rm d}\Xi_{1}(x)}{{\rm d}x}x\geq 0$ for $x>0$, indicating that the power-control policy proposed in [16] satisfies the condition of Theorem 8. There are two special cases to consider. When $a_{2}$ approaches 0, the power-control policy proposed in [16] simplifies to the water-filling policy, which has been discussed above. When $a_{2}$ approaches infinity, the power-control policy proposed in [16] becomes channel inversion policy, with which the rate remains constant over all channel states. In this case, EC is trivial, as the rate remains fixed.
@@
Although the approximation provided in Theorem 8 can be applied to policies satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$, its accuracy can be further enhanced through additional optimization tailored to the specific properties of each policy. For example, in the water-filling scheme, there exists a channel-gain threshold $h_{\rm th}$. If the channel gain falls below this threshold, the allocated power becomes 0, implying that the transmission rate remains constant in the FBL regime when $h[n]<h_{\rm th}$. [^9] The actual rate should be zero when $\gamma=0$. However, since we use the normal approximation as mentioned at the beginning of this section, we let the rate be equal to $\frac{\log_{2}N}{2}$ when $\gamma=0$, in line with Eq. (9). This setting can be easily adjusted by setting the rate to 0 when $\gamma=0$, which does not affect the accuracy of this approximation. Let $\xi(\theta)=e^{-\theta T\frac{\log_{2}(N)}{2}}\int_{0}^{h_{\rm th}}f(x){\rm d}x$. Based on the above analysis and Theorem 8, EC can be approximated using Eq. (78), which is presented at the bottom of this page. For those policies exhibiting truncated characteristics, Eq. (78) provides a more accurate approximation of their EC.
$\alpha_{S}(\theta)=-\frac{1}{\theta T}\ln\left(\epsilon+(1-\epsilon)\sqrt{\frac{2\pi}{\theta NT|r^{\prime\prime}(x^{*})|}}f(x^{*})e^{-\theta NTr(x^{*})}\left(1+O\left((\theta NT)^{-1}\right)\right)\right).$ (79)
For the low-latency system incorporating the retransmission mechanism, EC is discussed in [28], which differs from the analysis provided above where no retransmission mechanism is considered. In [28], a simple ARQ mechanism is assumed. Under this mechanism, the receiver sends a negative acknowledgment to request the retransmission of the message in case of an erroneous reception. In Corollary 4, we generalize the EC approximation to the low-latency communication system incorporating this ARQ mechanism.
Corollary 4. For a low-latency system utilizing the above simple ARQ mechanism and a channel-gain-based power allocation scheme $\Xi(x)$ that satisfies $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$, with channel distribution $f(\cdot)$, EC is approximated using Eq. (79), which is located at the bottom of this page. In Eq. (79), $r(x)$ is defined by Eq. (41), and $x^{*}$ is given by Eq. (43).
Proof:
See Appendix J. ∎
V Numerical Results
In this section, we perform numerical calculations to validate the theoretical results and conclusions presented in this paper. Moreover, the numerical results provide a more visual explanation of some characteristics of the statistical-delay performance in low-latency systems operating in the FBL regime.
First, we validate the conclusions regrading the high-SNR slopes of EC in the FBL regime, shown in Figs. 5, 5, and 6. In these figures, we set $N=512$, $\epsilon=10^{-5}$, and $T=1$. In Fig. 5, $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=1$. In Fig. 5, $f(x)$ is an Erlang distribution with $\kappa=2$ and $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=2$. In Fig. 6, we set $\Omega=1$. For the Rayleigh fading case, $\frac{1}{T\log_{2}e}\approx 0.6931$. From Fig. 5, we find that the slopes of $\Lambda(\gamma)$ for $\theta N\leq 0.6931$, i.e. $\theta\leq 0.0014$, are almost identical and close to 1, while the slopes of $\Lambda(\gamma)$ for $\theta>0.0014$ are smaller than 1. Moreover, for $\theta>0.0014$, a larger $\theta$ leads to a smaller slope of $\Lambda(\gamma)$. Specifically, the slope of $\Lambda(\gamma)$ for $\theta=0.004$ is $0.338$ in Fig. 5, while $\frac{1}{\theta NT\log_{2}e}$ with $\theta=0.004$ equals 0.3385. These results validate the accuracy of Corollary 1.
Next, we discuss the numerical results of the case in which $|h[n]|^{2}$ follows an Erlang distribution. In Fig. 5, $\frac{\kappa}{T\log_{2}e}\approx 1.3863$. We observe that for $\theta=0.002$, $\theta N=1.024$ is larger than $\frac{1}{T\log_{2}e}\approx 0.6931$ in Fig. 5, but smaller than $\frac{\kappa}{T\log_{2}e}\approx 1.3863$ in Fig. 5. Accordingly, the slope of $\Lambda(\gamma)$ for $\theta=0.002$ in Fig. 5 differs from the slope of ergodic capacity, while the slope of $\Lambda(\gamma)$ for $\theta=0.002$ is nearly identical to the slope of ergodic capacity in Fig. 5. This result highlights the differences in $s_{\infty}$ performance with and without spatial or frequency diversity in Rayleigh fading channels, further validating Corollary 1.